# The long way of a viscous vortex dipole

### v1

@@ 3.3. The induction step
$\mathcal{N}_{M+1}\,$ =$=\,\Bigl{\{}\Psi_{\mathrm{app}}-\Psi_{0}-\mathcal{T}_{\varepsilon}\Psi_{\mathrm{app}}+\frac{\varepsilon\xi_{1}}{2\pi}\,\zeta_{\mathrm{app}}\,,\,\Omega_{M+1}\Bigr{\}}$
+$+\Bigl{\{}\Psi_{M+1}+\frac{\xi_{1}}{2\pi}\,\zeta_{M}\,,\,\Omega_{\mathrm{app}}-\Omega_{0}+\varepsilon^{M+1}\Omega_{M+1}\Bigr{\}}-\Bigl{\{}\mathcal{T}_{\varepsilon}\Psi_{M+1}\,,\,\Omega_{\mathrm{app}}+\varepsilon^{M+1}\Omega_{M+1}\Bigr{\}}\,.$
@@
~$\tilde{\mathcal{H}}_{0}\,:=\,\mathcal{H}_{0}+\frac{\zeta_{M}^{E}}{2\pi}\,\partial_{2}\Omega_{0}\,,\qquad\tilde{\mathcal{H}}_{1}\,:=\,\mathcal{H}_{1}+\frac{\zeta_{M}^{NS}}{2\pi}\,\partial_{2}\Omega_{0}\,,$ (3.36)
@@
[…]
    Remark that ~$\tilde{\mathcal{H}}_{0}\in\mathrm{Ker}(\Lambda)^{\perp}$ because ~$\tilde{\mathcal{H}}_{0}(\xi)$ is an odd function of $\xi_{2}$, which implies that $\mathcal{P}_{0}\tilde{\mathcal{H}}_{0}=0$ and $\mathrm{m}_{1}[\tilde{\mathcal{H}}_{0}]=0$, and because $\mathrm{m}_{2}[\tilde{\mathcal{H}}_{0}]=0$ by our choice of $\zeta_{M}^{E}$. Note also that $\Omega_{M+1}^{E,1}$ is an even function of $\xi_{2}$, as asserted in Proposition 3.6.
@@ 3.4. Leading order correction to the vertical speed
We already know that $\Omega_{2}=\Omega_{2}^{E}+\delta\Omega_{2}^{NS}$ where $\Lambda\Omega_{2}^{E}+\mathcal{H}_{2}=0$ and $\Lambda\Omega_{2}^{NS}+(\mathcal{L}-1)\Omega_{2}^{E}=0$, see Section 3.2. According to (3.27) we have $-\mathcal{H}_{2}=b(r)\sin(2\theta)$ where $b(r)=r^{2}g(r)/(2\pi)$ with $g$ as in (A.8). In particular $\mathcal{H}_{2}\in\mathcal{Y}_{2}\cap\mathcal{Z}$, so that we can apply Lemma A.1 in Section A.3. We thus obtain the formulas $\Omega_{2}^{E}=w(r)\cos(2\theta)$, $\Psi_{2}^{E}=\varphi(r)\cos(2\theta)$, where
@@
[…]
@@
The error term $\mathcal{O}(\varepsilon^{5}+\delta^{2}\varepsilon)$ in (3.43) is probably not optimal. We believe that it can be improved to $\mathcal{O}(\varepsilon^{5}+\delta^{2}\varepsilon^{4})$, but this requires nontrivial modifications of our arguments. We also recall that $\varepsilon^{2}=\delta t/T_{\mathsf{adv}}$, where $T_{\mathsf{adv}}$ is defined in (1.12). If $t\gtrsim T_{\mathsf{adv}}$, then $\delta\lesssim\varepsilon^{2}$ and the error term in (3.43) can be replaced by $\mathcal{O}(\varepsilon^{5})$.
@@ 3.5. Functional relationship in the inviscid case
where the last term in the right-hand side is an irrelevant constant that is included for convenience. The remainder $\mathcal{R}_{M}^{E}=\bigl{\{}\Phi_{\mathrm{app}}^{E}\,,\,\Omega_{\mathrm{app}}^{E}\bigr{\}}$ would vanish identically if we had a functional relationship of the form $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})=0$ for some (smooth) function $F:\mathbb{R}\to\mathbb{R}$. In reality, since $\mathcal{R}_{M}^{E}=\mathcal{O}(\varepsilon^{M+1})$, the best we can hope for is an approximate functional relationship, which holds up to corrections of order $\mathcal{O}(\varepsilon^{M+1})$.
@@
$A(\xi)\,:=\,F_{0}^{\prime}\bigl{(}\Omega_{0}(\xi)\bigr{)}\,=\,-\frac{\nabla\Psi_{0}(\xi)}{\nabla\Omega_{0}(\xi)}\,=\,\frac{4}{|\xi|^{2}}\Bigl{(}e^{|\xi|^{2}/4}-1\Bigr{)}\,.$ (3.51)
@@
The level lines of the second order approximation computed above are shown in Fig. 1 (ignoring the $\varepsilon^{3}$ corrections). We now look for a relationship of the form $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})=\mathcal{O}_{\mathcal{S}_{*}}(\varepsilon^{3})$, where $F=F_{0}+\varepsilon^{2}F_{2}$ for some $F_{2}:(0,+\infty)\to\mathbb{R}$. This is problematic, however, because it is not true that $\Omega_{\mathrm{app}}^{E}(\xi)>0$ for all $\xi\in\mathbb{R}^{2}$. As is explained in Remark 3.21 below, this difficulty is avoided if we only require that the second order Taylor polynomial in $\varepsilon$ of the quantity $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})$ vanishes. This gives the relation
[…]
@@
[…]
@@
To simplify the notation, we drop the superscript “$E$” and the subscript “$\mathrm{app}$” everywhere. We proceed by induction over the integer $M$. Assume that, for some $M\geq 2$, we have constructed $\Phi\in\mathcal{S}_{*}(\mathbb{R}^{2})$, $\Omega\in\mathcal{Z}$, and $F\in\mathcal{K}$ of the form
@@
Finally using (3.56), (3.57) we find
$\Pi_{M+1}\bigl{(}\tilde{\Phi}+\tilde{F}(\tilde{\Omega})\bigr{)}\,$ $=\,\Pi_{M+1}\bigl{(}\Phi+F(\Omega)\bigr{)}+\Pi_{M+1}\bigl{(}\tilde{\Phi}-\Phi\bigr{)}+\Pi_{M+1}\bigl{(}\tilde{F}(\tilde{\Omega})-F(\Omega)\bigr{)}$
@@
$\bigl{\{}F^{\prime}(\Omega_{0})\Omega_{M+1}\,,\,\Omega_{0}\bigr{\}}\,=\,\bigl{\{}\Omega_{M+1}\,,\,F(\Omega_{0})\bigr{\}}\,=\,\bigl{\{}\Omega_{M+1}\,,\,-\Phi_{0}\bigr{\}}\,=\,\bigl{\{}\Phi_{0}\,,\,\Omega_{M+1}\bigr{\}}\,.$
@@
[…]
@@
[…]
@@ 4. Correction to the approximate solution
$Z_{2}^{\prime}(t)\,=\,\frac{\Gamma}{2\pi d}\bigl{(}\zeta_{\mathrm{app}}(t)+\zeta(t)\bigr{)}\,,\qquad$ (4.2)
@@
+$+\Bigl{\{}\varphi-\mathcal{T}_{\varepsilon}\varphi\,,w\Bigr{\}}+\frac{1}{\delta^{2}}\,\mathcal{R}_{M}+\frac{\varepsilon\,\zeta}{2\pi\delta^{2}}\Bigl{\{}\xi_{1},\Omega_{\mathrm{app}}+\delta w\Bigr{\}}\,.$ (4.5)
@@ 4.1. Main difficulties and Arnold’s variational principle
+$+\bigl{\{}\varphi-\mathcal{T}_{\varepsilon}\varphi\,,w\bigr{\}}+\frac{1}{\delta^{2}}\,\mathcal{R}_{M}+\frac{\varepsilon\,\zeta}{2\pi\delta^{2}}\bigl{\{}\xi_{1},\Omega_{\mathrm{app}}+\delta w\bigr{\}}\,.$ (4.11)
@@
′′$\mathcal{E}^{\prime\prime}(\omega_{*})[\omega,\omega]\,=\,\int F^{\prime}(\omega_{*})\omega^{2}\,{\rm d}x-\int\psi\omega\,{\rm d}x\,.$ (4.13)
A key observation is that this quadratic function of $\omega$ is invariant under the evolution defined by the linearized Euler equation at $\omega_{*}$. Indeed, if $\partial_{t}\omega=-\{\psi_{*},\omega\}-\{\psi,\omega_{*}\}$, a direct calculation shows that ′′$\mathcal{E}^{\prime\prime}(\omega_{*})[\omega,\omega]$ does not vary in time. As a consequence, if the quadratic form (4.13) has a definite sign, it can be used to prove the stability of the steady state $\omega_{*}$, not only for the Euler equation but for related systems as well, see [13] for an application to the stability of vortices in the 2D Navier-Stokes equations.
@@
[…]
@@
[…]
@@ 4.2. The weighted space and its properties
[…]
@@
$\mathrm{I}_{\varepsilon}\,$ =$=\,\bigl{\{}\xi\in\mathbb{R}^{2}\,:\,|\xi|<2\varepsilon^{-\sigma_{1}}\,,~{}F^{\prime}\bigl{(}\Omega_{\mathrm{app}}^{E}(\xi)\bigr{)}<\exp(\varepsilon^{-2\sigma_{1}}/4)\bigr{\}}\,,\hskip 22.76228pt$ (Inner)
$\mathrm{II}_{\varepsilon}\,$ =$=\,\bigl{\{}\xi\in\mathbb{R}^{2}\,:\,\xi\notin\mathrm{I}_{\varepsilon}\,,~{}|\xi|\leq\varepsilon^{-\sigma_{2}}\bigr{\}}\,,$ (Intermediate)
$\mathrm{III}_{\varepsilon}\,$ =$=\,\bigl{\{}\xi\in\mathbb{R}^{2}\,:\,|\xi|>\varepsilon^{-\sigma_{2}}\bigr{\}}\,,$ (Outer)
@@
Figure 2. A schematic representation of the graph of the weight function $W_{\varepsilon}$ defined in (4.22). In the inner region $\mathrm{I}_{\varepsilon}$, the weight is close for $\varepsilon>0$ small to the radially symmetric function $W_{0}(\xi)=4|\xi|^{-2}\bigl{(}e^{|\xi|^{2}/4}-1\bigr{)}$. It then takes constant values in the intermediate region $\mathrm{II}_{\varepsilon}$, and grows like $\exp(|\xi|^{2\gamma}/4)$ in the outer region $\mathrm{III}_{\varepsilon}$. The dashed lines illustrate the bounds (4.25), where the constants $C_{1},C_{2}$ are independent of $\varepsilon$.
@@
We collect elementary properties of the inner region $\mathrm{I}_{\varepsilon}$ and of the weight function $W_{\varepsilon}$ in the following lemma, which is the analogue of [14, Lemmas 4.2 & 4.3].
@@
We next derive useful estimates for the stream function and the velocity field in terms of the vorticity in our function space $\mathcal{X}_{\varepsilon}$.
@@
Let $w\in\mathcal{X}_{\varepsilon}$ and $\varphi=\Delta^{-1}w$ as in (2.6). Then, for all $2<q<\infty$ there exists a constant $C>0$ such that
@@
Moreover
@@
Proof.
In view of (2.6) we have, for all $\xi\in\mathbb{R}^{2}$,
$|\varphi(\xi)|\,\lesssim\,\int_{\mathbb{R}^{2}}\bigl{|}\log|\xi-\eta|\bigr{|}\,|w(\eta)|\,{\rm d}\eta\,,\quad|\nabla\varphi(\xi)|\,\lesssim\,\int_{\mathbb{R}^{2}}\frac{1}{|\xi-\eta|}\,|w(\eta)|\,{\rm d}\eta\,.$ (4.33)
The bound (4.31) on $\nabla\varphi$ is a direct consequence of (4.33) and the Hardy-Littlewood-Sobolev inequality, see [15, Lemma 2.1]. More precisely, given $1<p<2<q<\infty$ with $21/p=1/q+1/$ 2, the HLS inequality shows that $\left\lVert\nabla\varphi\right\rVert_{L^{q}}\lesssim\left\lVert w\right\rVert_{L^{p}}\lesssim\left\lVert w\right\rVert_{\mathcal{X}_{\varepsilon}}$, where in the last step we used the fact that $\mathcal{X}_{\varepsilon}\hookrightarrow L^{r}(\mathbb{R}^{2})$ for any $r\in[1,2]$.
To conclude the proof of (4.31), we use the first inequality in (4.33) together with the crude bounds $|\log|\xi-\eta||\leq|\xi-\eta|^{-1/2}$ if $|\xi-\eta|\leq 1$, and $\log|\xi-\eta|\leq\log(1+|\xi|)+\log(1+|\eta|)$ if $|\xi-\eta|\geq 1$. We thus find
$|\varphi(\xi)|\,\lesssim\,\int_{\{|\xi-\eta|\leq 1\}}\frac{1}{|\xi-\eta|^{1/2}}|w(\eta)|\,{\rm d}\eta+\log(1+|\xi|)\int_{\{|\xi-\eta|\geq 1\}}\log(2+|\eta|)|w(\eta)|\,{\rm d}\eta\,,$
and applying Hölder’s inequality to the first integral we arrive at $|\varphi(\xi)|\leq C\log(2+|\xi|)\|w\|_{\mathcal{X}_{\varepsilon}}$. This implies in particular that $(1+|\cdot|)^{-1}\varphi\in L^{q}(\mathbb{R}^{2})$ for all $q>2$, as asserted in (4.31).
Finally, to prove (4.32), we deduce from the second inequality in (4.33) that
The last integrand is equal to $|\xi-\eta|^{-1}\bigl{(}(1+|\eta|)|w(\eta)|^{1/2}\bigr{)}|w(\eta)|^{1/2}$, so we can apply the trilinear Hölder inequality with exponents $8/5$, $4$, $8$ to obtain
where in the last step we used the embedding $H^{1}(\mathbb{R}^{2})\hookrightarrow L^{4}(\mathbb{R}^{2})$. This gives (4.32). ∎
Our second estimate focuses on the translated stream function $\mathcal{T}_{\varepsilon}\varphi$ near the origin.
@@
Assume that $w\in\mathcal{X}_{\varepsilon}$ with $\mathrm{M}[w]=\mathrm{m}_{1}[w]=\mathrm{m}_{2}[w]=0$ and let $\varphi=\Delta^{-1}w$ as in (2.6). Then, for any $q>2$ there exists a constant $C>0$ such that
$\varepsilon\bigl{\|}\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-2}\mathcal{T}_{\varepsilon}\varphi\bigr{\|}_{L^{q}}+\bigl{\|}\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-3}\nabla\mathcal{T}_{\varepsilon}\varphi\bigl{\|}_{L^{q}}\,\leq\,C\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,.$ (4.34)
Moreover $\|(1+|\cdot|)\varphi\|_{L^{\infty}}\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$.
The proof combines the expansion established in Lemma 3.8 with standard estimates exploiting the coercivity properties of the weight $W_{\varepsilon}$. The argument is somewhat technical, so we postpone it to Appendix A.4.

### v2

@@ 3.3. The induction step
$\mathcal{N}_{M+1}\,$ $=\,\Bigl\{\Psi_{\mathrm{app}}-\Psi_{0}-\mathcal{T}_{\varepsilon}\Psi_{\mathrm{app}}+\frac{\varepsilon\xi_{1}}{2\pi}\,\zeta_{\mathrm{app}}\,,\,\Omega_{M+1}\Bigr\}$
$+\Bigl\{\Psi_{M+1}+\frac{\xi_{1}}{2\pi}\,\zeta_{M}\,,\,\Omega_{\mathrm{app}}-\Omega_{0}+\varepsilon^{M+1}\Omega_{M+1}\Bigr\}-\Bigl\{\mathcal{T}_{\varepsilon}\Psi_{M+1}\,,\,\Omega_{\mathrm{app}}+\varepsilon^{M+1}\Omega_{M+1}\Bigr\}\,.$
@@
$\tilde{\mathcal{H}}_{0}\,:=\,\mathcal{H}_{0}+\frac{\zeta_{M}^{E}}{2\pi}\,\partial_{2}\Omega_{0}\,,\qquad\tilde{\mathcal{H}}_{1}\,:=\,\mathcal{H}_{1}+\frac{\zeta_{M}^{NS}}{2\pi}\,\partial_{2}\Omega_{0}\,$, (3.36)
@@
    $\Lambda\Omega_{M+1}^{E,1}+\tilde{\mathcal{H}}_{0}\,=\,0\,.$ (3.39)
    Remark that $\tilde{\mathcal{H}}_{0}\in\mathrm{Ker}(\Lambda)^{\perp}$ because $\tilde{\mathcal{H}}_{0}(\xi)$ is an odd function of $\xi_{2}$, which implies that $\mathcal{P}_{0}\tilde{\mathcal{H}}_{0}=0$ and $\mathrm{m}_{1}[\tilde{\mathcal{H}}_{0}]=0$, and because $\mathrm{m}_{2}[\tilde{\mathcal{H}}_{0}]=0$ by our choice of $\zeta_{M}^{E}$. Note also that $\Omega_{M+1}^{E,1}$ is an even function of $\xi_{2}$, as asserted in Proposition 3.6.
@@ 3.4. Leading order correction to the vertical speed
We already know that $\Omega_{2}=\Omega_{2}^{E}+\delta\Omega_{2}^{NS}$ where $\Lambda\Omega_{2}^{E}+\mathcal{H}_{2}=0$ and $\Lambda\Omega_{2}^{NS}+(\mathcal{L}-1)\Omega_{2}^{E}=0$, see Section 3.2. According to (3.27) we have $\mathcal{H}_{2}=b(r)\sin(2\theta)$ where $b(r)=r^{2}g(r)/(2\pi)$ with $g$ as in (A.8). In particular $\mathcal{H}_{2}\in\mathcal{Y}_{2}\cap\mathcal{Z}$, so that we can apply Lemma A.1 in Section A.3. We thus obtain the formulas $\Omega_{2}^{E}=w(r)\cos(2\theta)$, $\Psi_{2}^{E}=\varphi(r)\cos(2\theta)$, where
@@
$\alpha\,=\,\frac{1}{\pi}\int_{\mathbb{R}^{2}}\bigl(\xi_{2}^{2}-\xi_{1}^{2}\bigr)\Omega_{2}(\xi)\,{\rm d}\xi\,=\,\int_{0}^{\infty}r^{3}\mathsf{w}_{2}(r)\,{\rm d}r\,\approx\,22.24\,.$ (3.44)
@@
The error term $\mathcal{O}(\varepsilon^{5}+\delta^{2}\varepsilon)$ in (3.43) is probably not optimal, and may be improved using the techniques developed in [10]. We just recall here that $\varepsilon^{2}=\delta t/T_{\mathsf{adv}}$, where $T_{\mathsf{adv}}$ is defined in (1.12). Thus, if $t\gtrsim T_{\mathsf{adv}}$, then $\delta\lesssim\varepsilon^{2}$ and the error term in (3.43) can be replaced by $\mathcal{O}(\varepsilon^{5})$.
@@ 3.5. Functional relationship in the inviscid case
where the last term on the right-hand side is a constant included in order to cancel an irrelevant $\log(1/\varepsilon)$-order in the expansion (3.52) below. The remainder $\mathcal{R}_{M}^{E}=\bigl\{\Phi_{\mathrm{app}}^{E}\,,\,\Omega_{\mathrm{app}}^{E}\bigr\}$ would vanish identically if we had a functional relationship of the form $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})=0$ for some (smooth) function $F:\mathbb{R}\to\mathbb{R}$. In reality, since $\mathcal{R}_{M}^{E}=\mathcal{O}(\varepsilon^{M+1})$, the best we can hope for is an approximate functional relationship, which holds up to corrections of order $\mathcal{O}(\varepsilon^{M+1})$.
@@
$A(\xi)\,:=\,F_{0}^{\prime}\bigl(\Omega_{0}(\xi)\bigr)\,=\,-\frac{\nabla\Psi_{0}(\xi)}{\nabla\Omega_{0}(\xi)}\,=\,\frac{4}{|\xi|^{2}}\Bigl({\rm e}^{|\xi|^{2}/4}-1\Bigr)\,.$ (3.51)
@@
The level lines of the second order approximation (3.52) are shown in Fig. 1 (ignoring the $\varepsilon^{3}$ corrections). We now look for a relationship of the form $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})=\mathcal{O}_{\mathcal{S}_{*}}(\varepsilon^{3})$, where $F=F_{0}+\varepsilon^{2}F_{2}$ for some $F_{2}:(0,+\infty)\to\mathbb{R}$. This is problematic, however, because it is not true that $\Omega_{\mathrm{app}}^{E}(\xi)>0$ for all $\xi\in\mathbb{R}^{2}$. As is explained in Remark 3.22 below, this difficulty is avoided if we only require that the second order Taylor polynomial in $\varepsilon$ of the quantity $\Phi_{\mathrm{app}}^{E}+F(\Omega_{\mathrm{app}}^{E})$ vanishes. This gives the relation
$\Psi_{2}^{E}+\frac{1}{4\pi}\,\bigl(\xi_{1}^{2}-\xi_{2}^{2}\bigr)+F_{0}^{\prime}(\Omega_{0})\Omega_{2}^{E}+F_{2}(\Omega_{0})\,=\,0\,$, (3.53)
@@
$2\nabla\mathcal{F}_{k}(\xi)+\xi\bigl(\mathcal{F}_{k+1}(\xi)+k\mathcal{F}_{k}(\xi)\bigr)\,=\,0\,,\qquad\xi\in\mathbb{R}^{2}\,.$ (3.54)
@@
To simplify the notation, we drop the superscript “EE” and the subscript $\mathrm{app}$” everywhere. We proceed by induction over the integer $M$. Assume that, for some $M\geq 2$, we have constructed $\Phi\in\mathcal{S}_{*}(\mathbb{R}^{2})$, $\Omega\in\mathcal{Z}$, and $F\in\mathcal{K}$ of the form
@@
$\bigl\{\mathcal{H}_{M+1}\,,\,\Omega_{0}\bigr\}+\bigl\{\Phi_{M+1}\,,\,\Omega_{0}\bigr\}+\bigl\{\Phi_{0}\,,\,\Omega_{M+1}\bigr\}\,=\,0\,.$ (3.58)
Finally using (3.56), (3.57) we find
@@
$\bigl\{F_{0}^{\prime}(\Omega_{0})\Omega_{M+1}\,,\,\Omega_{0}\bigr\}\,=\,\bigl\{\Omega_{M+1}\,,\,F_{0}(\Omega_{0})\bigr\}\,=\,\bigl\{\Omega_{M+1}\,,\,-\Phi_{0}\bigr\}\,=\,\bigl\{\Phi_{0}\,,\,\Omega_{M+1}\bigr\}\,.$
@@
$\frac{1}{2}\,\Omega_{0}(\xi)\,\leq\,\Omega_{\mathrm{app}}^{E}(\xi)\,\leq\,2\Omega_{0}(\xi)\,,\qquad|\xi|\,\leq\,2\varepsilon^{-\sigma_{1}}\,$, (3.61)
@@
$\Theta(\varepsilon,\xi)\,=\,\frac{\varepsilon^{M+1}}{M!}\int_{0}^{1}(1-\tau)^{M}\partial_{\varepsilon}^{M+1}\Theta(\tau\varepsilon,\xi)\,{\rm d}\tau\,,\qquad|\xi|\,\leq\,2\varepsilon^{-\sigma_{1}}\,.$ (3.62)
@@ 4. Correction to the approximate solution
$Z_{2}^{\prime}(t)\,=\,\frac{\Gamma}{2\pi d}\bigl(\zeta_{\mathrm{app}}(t)+\zeta(t)\bigr)\,$$\qquad$ (4.2)
@@
$+\Bigl\{\varphi-\mathcal{T}_{\varepsilon}\varphi\,,w\Bigr\}+\frac{1}{\delta^{2}}\,\mathcal{R}_{M}+\frac{\varepsilon\,\zeta}{2\pi\delta^{2}}\Bigl\{\xi_{1},\Omega_{\mathrm{app}}+\delta w\Bigr\}\,.$ (4.5)
@@ 4.1. Main difficulties and Arnold’s variational principle
$+\bigl\{\varphi-\mathcal{T}_{\varepsilon}\varphi\,,w\bigr\}+\frac{1}{\delta^{2}}\,\mathcal{R}_{M}+\frac{\varepsilon\,\zeta}{2\pi\delta^{2}}\bigl\{\xi_{1},\Omega_{\mathrm{app}}+\delta w\bigr\}\,.$ (4.11)
@@
$\mathcal{E}^{\prime\prime}(\omega_{*})[\omega,\omega]\,=\,\int F^{\prime}(\omega_{*})\omega^{2}\,{\rm d}x-\int\psi\omega\,{\rm d}x\,.$ (4.13)
A key observation is that this quadratic function of $\omega$ is invariant under the evolution defined by the linearized Euler equation at $\omega_{*}$. Indeed, if $\partial_{t}\omega=-\{\psi_{*},\omega\}-\{\psi,\omega_{*}\}$, a direct calculation shows that $\mathcal{E}^{\prime\prime}(\omega_{*})[\omega,\omega]$ does not vary in time. As a consequence, if the quadratic form (4.13) has a definite sign, it can be used to prove the stability of the steady state $\omega_{*}$, not only for the Euler equation but for related systems as well, see [14] for an application to the stability of vortices in the 2D Navier-Stokes equations.
@@
= $=\,0+\int w\bigl\{\varphi-\mathcal{T}_{\varepsilon}\varphi+\zeta\xi_{1}\,,F(\Omega_{\mathrm{app}}^{E})\bigr\}\,{\rm d}\xi\,$,
@@
= $=\,0+\int\bigl(\varphi-\mathcal{T}_{\varepsilon}\varphi\bigr)\Bigl(\bigl\{w\,,F(\Omega_{\mathrm{app}}^{E})\bigr\}+\zeta\bigl\{\xi_{1}\,,\Omega_{\mathrm{app}}^{E}\bigr\}\Bigr)\,{\rm d}\xi\,.$
@@ 4.2. The weighted space and its properties
$0\,<\,\sigma_{1}\,<\,\frac{1}{2}\,,\qquad\sigma_{2}\,>\,1\,,\qquad\gamma\,=\,\frac{\sigma_{1}}{\sigma_{2}}\,<\,\frac{1}{2}\,.$ (4.19)
@@
[…]
$\mathrm{II}_{\varepsilon}\,$ $=\,\bigl\{\xi\in\mathbb{R}^{2}\,:\,\xi\notin\mathrm{I}_{\varepsilon}\,,~|\xi|\leq\varepsilon^{-\sigma_{2}}\bigr\}\,$, (Intermediate)
$\mathrm{III}_{\varepsilon}\,$ $=\,\bigl\{\xi\in\mathbb{R}^{2}\,:\,|\xi|>\varepsilon^{-\sigma_{2}}\bigr\}\,$, (Outer)
@@
Figure 2. A schematic representation of the graph of the weight function $W_{\varepsilon}$ defined in (4.22). In the inner region $\mathrm{I}_{\varepsilon}$, the weight is close for $\varepsilon>0$ small to the radially symmetric function $W_{0}(\xi)=4|\xi|^{-2}\bigl({\rm e}^{|\xi|^{2}/4}-1\bigr)$. It then takes constant values in the intermediate region $\mathrm{II}_{\varepsilon}$, and grows like $\exp(|\xi|^{2\gamma}/4)$ in the outer region $\mathrm{III}_{\varepsilon}$. The dashed lines illustrate the bounds (4.25), where the constants $C_{1},C_{2}$ are independent of $\varepsilon$.
@@
We collect elementary properties of the inner region $\mathrm{I}_{\varepsilon}$ and of the weight function $W_{\varepsilon}$ in the following lemma, which is the analogue of [15, Lemmas 4.2 & 4.3].
@@
Assuming that the vorticity lies in the function space $\mathcal{X}_{\varepsilon}$ for some $\varepsilon>0$, we next recall a few classical estimates for the stream function and the velocity field. For the reader’s convenience, the proof of the following statements is reproduced in Section A.4.
@@
If $w\in\mathcal{X}_{\varepsilon}$ and $\varphi=\Delta^{-1}w$ in the sense of (2.6), there exists a constant $C>0$ such that $|\varphi(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}\log(2+|\xi|)$ for all $\xi\in\mathbb{R}^{2}$. Moreover, for all $q\in(2,\infty)$, there holds
@@
If in addition $\nabla w_{\varepsilon}\in\mathcal{X}_{\varepsilon}$, then
@@
The conclusions of Lemma 4.5 can be considerably strengthened if the mean and the first order moments of $w$ are assumed to vanish.
@@
Assume that $w\in\mathcal{X}_{\varepsilon}$ satisfies $\mathrm{M}[w]=\mathrm{m}_{1}[w]=\mathrm{m}_{2}[w]=0$, and let $\varphi=\Delta^{-1}w$ as in (2.6). Then for all $q>2$ there exists a constant $C>0$ such that
$\|(1+|\cdot|)^{2}\varphi\|_{L^{\infty}}+\|(1+|\cdot|)^{2}\nabla\varphi\|_{L^{q}}\,\leq\,C\left\lVert w\right\rVert_{\mathcal{X}_{\varepsilon}}.$ (4.33)
If in addition $\nabla w_{\varepsilon}\in\mathcal{X}_{\varepsilon}$, then
$\|(1+|\cdot|)^{3}\nabla\varphi\|_{L^{\infty}}\,\leq\,C\bigl(\left\lVert\nabla w\right\rVert_{\mathcal{X}_{\varepsilon}}^{1/2}+\left\lVert w\right\rVert_{\mathcal{X}_{\varepsilon}}^{1/2}\bigr)\left\lVert w\right\rVert_{\mathcal{X}_{\varepsilon}}^{1/2}\,.$ (4.34)
The bounds (4.33), (4.34) will be used in particular to estimate the translated stream function $\mathcal{T}_{\varepsilon}\varphi$ in the inner region near the origin.
Corollary 4.7.
Under the assumptions of Lemma 4.6 we have, for any $q\in(2,\infty)$,
$\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}\mathcal{T}_{\varepsilon}\varphi\bigr\|_{L^{\infty}}\,\leq\,C\varepsilon^{2}\|w\|_{\mathcal{X}_{\varepsilon}}\,,\qquad\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(1+|\cdot|)^{-1}\nabla\mathcal{T}_{\varepsilon}\varphi\bigl\|_{L^{q}}\,\leq\,C\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,.$ (4.35)